# HyperMOOC: Augmenting MOOC Videos with Concept-based Embedded Visualizations

### v1

@@ 1. Introduction
With the rise of distance learning and online education platforms, Massive Open Online Courses (MOOCs) have continued to attract new learners over the past decade. However, learners often find it confusing to watch MOOC videos without the aid of additional knowledge(Milligan and Griffin, 2016) due to their lack of sufficient knowledge to understand the complex arrangement of the course and the diverse logical relationships of concepts - the fundamental ideas, theories, and principles that form the core of the course content. Existing methods of providing additional information, such as navigation menus (Zhang et al., 2019; Zhou et al., 2024) and course overviews (Rahman et al., 2023; Schwab et al., 2016), usually fail to address their confusion adequately. These methods either fail to provide on-demand assistance or distract learners’ attention by displaying concept flows in separate views. Thus, finding a way to seamlessly access additional data while watching a MOOC video is crucial for effectively learning and understanding a course.
@@
Embedded visualizations present a promising opportunity to access additional data by directly displaying them within actual scenes, without diverting user’s attention from MOOC videos. Various commercial products (Clippers, 2020; Vizrt, 2022) and research systems (Chen et al., 2023, 2021) have leveraged embedded visualizations to augment videos. However, these systems either do not focus on the educational domain or only utilize basic non-interactive text labels and progress bars to display data. In addition, hypervideo (Chambel et al., 2006), as an emerging interactive video technology, allows users to access additional information, media, or video segments on-demand by embedding hyperlinks within the video. As a means of learning in MOOCs, hypervideo has proven to have a positive impact on promoting knowledge acquisition and learner engagement (Milligan and Griffin, 2016). However, to the best of our knowledge, there is still no work focused on hypervideo-based augmentation for MOOC videos. Similar to all learning content analysis systems, the challenges of analysis mainly come from the Object (learning content) and the Subject (learner) of MOOC video learning: 1) learning content-level. What kinds of data can be used to augment MOOC videos (Q1)?, What visual effects are used to augment these data (Q2)?; 2) learner-level. How to dynamically augment data based on learners’ viewing behavior (Q3)?, How to augment data without increasing the cognitive load on learners while promoting their understanding of the course (Q4)?.
@@
In this work, we aim to fill this gap by developing interactive embedded visualizations to help learners watch MOOC videos. To understand learners’ specific practices, pain points, and existing solutions for watching MOOC videos, we systematically analyzed the previous literature, learners’ feedback from real learning scenarios, and insights gained from multiple rounds of expert interviews. Given that learning is primarily concept-driven (Latham, 2020), our analysis resulted in a concept-based design space to address the four aforementioned questions. The design space describes how to effectively augment MOOC videos in terms of both learning content and learner levels. The four dimensions of the design space (i.e., data type, visual effects, behavior level, and cognitive level) provide guidance for augmentation of MOOC videos, offering new perspectives for facilitating learning in the online education domain.
@@
Guided by the design space, we designed and implemented HyperMOOC, a MOOC video augmentation tool that enhances learners’ engagement and understanding of MOOC videos (Fig. 1-c). HyperMOOC uses raw MOOC videos as input (Fig. 1-a), extracts the elements at each element level (Sec. 3.3) through deep learning (DL) models, and implements a hyperlink-based interaction method to embed visualizations into MOOC video content (Fig. 1-b). Our interaction flow is based on our design space and typical MOOC video viewing behaviors, where learners can access different modes of visual augmentations and course understanding in three stages, i.e., “Play”, “Focused”, and “Paused”. These stages correspond to the common viewing patterns of MOOC learners, i.e., continuous playback, focused attention on specific content, and pausing for deeper engagement. Using a keyboard-mouse interaction, learners can seamlessly transition between these stages, thus balancing explanatory and exploratory aspects within MOOC video learning. In the paused stage, the tool displays the complete visual designs. Guided by the multi-stage methodology of knowledge acquisition (Mccarthy, 1996), we implemented a set of concept-based glyphs to visually encode the various types of basic data elements (figures & tables, equations & codes, examples & tests) for the concept description. A user study with 36 learners demonstrated the overall usability of HyperMOOC. Based on the three stages, we compared the learning experience between raw video (RAW), explanatory augmentation (AUG), and explanatory + exploratory augmentation (FULL). Participants agreed that the FULL version of HyperMOOC can more effectively promote the course understanding than the other two baselines. We further collected feedback on HyperMOOC from two educational psychologists, and discussed the limitations of the tool and inspiration for future research. The corpus, created videos, and other materials can be found in https://hypermooc.github.io/HyperMOOC/. In summary, the contributions of this paper can be summarized as follows:
@@ 2. Related Work
Object-level analysis focuses on indexing, retrieving, and browsing key concepts of online courses. To mitigate the overlooking of essential teaching content in low-quality courses, Haubold et al.(Haubold, 2004) delved into the analysis of extracting index words from audio transcriptions and mapped them as transcript index maps. Adcock et al.(Adcock et al., 2010) identified slides in lecture videos, and applied OCR and keyword recognition to generate a searchable list of terms, enhancing the user experience in searching and browsing lectures. Additionally, Albahr et al.(Albahr et al., 2019) proposed an unsupervised method called SemKeyphrase for extracting keywords from MOOC videos, which reduced the difficulty of obtaining key concepts.
@@
Event-level analysis focuses on identifying patterns between concepts and analyzing their interrelationship during concept transitions. Concept mapping plays a crucial role in knowledge representation, self-learning, reflection, and assessment(Chang et al., 2001; Oliver, 2009). Cañas et al.(Cañas and Novak, 2014) combined CmapTools and the Internet to enrich learning experiences. However, crafting high-quality concept maps typically demands considerable effort from learners. Therefore, Zhang et al.(Zhang et al., 2019) introduced ScaffoMapping to assist video learners in concept mapping and provided auxiliary functions to enhance the quality of concept maps. Additionally, Liu et al.(Liu et al., 2018) proposed ConceptScape to generate expert-level concept maps for lecture videos.
@@
Conclusion-level analysis focuses on providing course summaries to help learners grasp core concepts and content. For example, Haubold et al.(Haubold and Kender, 2005) segmented lecture videos and incorporated visual hints to facilitate video content browsing and searching. Furthermore, Zhao et al.(Zhao et al., 2016) provided a comprehensive content overview of biomedical videos through keyframes and important semantic information extracted from the videos. The reorganization of lecture videos into multiple sub-videos will make it easier for learners to index the content of interest. For example, Rahman et al.(Rahman et al., 2023) proposed an automated segmentation method for lecture videos, improving navigation and search efficiency.
Our contribution is proposing a concept-based augmented MOOC video design space that combines three levels of analysis. This design space is developed through discussions with experts and is based on the learning content. It aids learners in exploring and visualizing complex learning content, thereby enhancing their understanding and engagement with course materials.
@@
Online Learning Video Visualizations. Numerous studies have explored diverse methods to visualize online learning videos, aiming to enhance educators’ and learners’ understanding and engagement with educational content, leading to a variety of visualization needs.
To track student progress and activity in a video, it is crucial to perform a visual analysis of the clickstream. For example, Chen et al.(Chen et al., 2015) introduced PeakVizor, utilizing clock-like glyphs to explore learner peak relationships in online videos. However, clickstream analysis may suffer from data inaccuracies and can’t fully capture students’ learning conditions. Hence, He et al.(He et al., 2018) proposed VUSphere, a visual analytic system employing a spherical layout to analyze video, course, and student elements on a sphere. Chen et al.(Chen et al., 2018) introduced ViSeq, a visualization learning sequence analysis system utilizing bidirectional chord diagrams to understand the correlation between learning sequences and performance.
In terms of visualizing learning content, several works visually present concepts to make learning content more engaging. For example, Zhao et al.(Zhao et al., 2017) presented keyframe summaries in a comic book style, aiding users in intuitively grasping significant visual content. Zhou et al.(Zhou et al., 2024) proposed ConceptThread, presenting course concepts to learners in a thread view, facilitating understanding of concepts and their relationships. While ConceptThread aims to provide an overview of course content structure, our work has a different purpose. HyperMOOC focuses on augmenting the real-time video watching experience. We achieve this by directly integrating visualizations into the MOOC videos, allowing learners to maintain awareness of the knowledge context without interrupting their viewing flow. This approach addresses the limitation of frequent eye movements and attention shifts that often occur in external visualization methods. To the best of our knowledge, we are the first to apply embedded visualization to augment MOOC videos, offering a more seamless and cognitively continuous learning experience.
@@
The dynamic nature of hypervideo transforms static information into engaging and motion-based content, enhancing understanding and knowledge transfer in educational environments. For instance, Chambel et al.(Chambel et al., 2006) demonstrated that hypervideo could visualize static dimensions as dynamic ones, thus promoting comprehension. In MOOCs, this dynamism is particularly beneficial, allowing viewers to access additional information about video clips through hyperlinks.
Control features in hypervideo systems empower users to navigate and interact with video content more flexibly. HyperMeeting(Girgensohn et al., 2015) exemplifies this by offering extensive control over video playback, allowing participants to initiate playback for all current participants, navigate between previously recorded meetings, and pause or skip recorded video streams.
Hyperlink is central to hypervideo, enabling the creation of interconnected video content and additional information. The inter-video navigation capability is often complemented by annotation and discussion features, as seen in TrACE(Dorn et al., 2015), which allows viewers to annotate and embed discussions at specific locations and times in videos. In educational applications, Tiellet et al.(Tiellet et al., 2010) created MultiLinks, which allows viewing surgeries organized in steps, linked to corresponding times in the hypervideo, providing effective learning support in veterinary surgery education.
HyperMOOC builds upon traditional hypervideo systems by integrating features such as conceptual embedded visualization, multi-symbol design, multi-stage interaction, cognitive path visualization, and hyperlink-based course navigation. This approach not only enhances learners’ understanding of MOOC video content and grasp of knowledge frameworks but also improves learning efficiency and experience.
@@ 3.1. Methodology
Literature Review. We employed both search-driven and reference-driven methods for the literature review. Specifically, for the search-driven method, we used the search keywords “online learning”, “education videos”, “MOOC videos”, “learning content analysis”, and “e-learning analysis” to identify related research papers. We conducted our research paper search using IEEE Explorer, ACM Digital Library, and Google Scholar, focusing primarily on high-impact conferences (such as IEEE VIS, ACM CHI, ACM L@S, and ACM LAK) and journals (including IEEE TVCG, Computer Graphics & Applications, and Computer & Education). After our first round of search, we recursively searched for papers in the references. Next, we browsed each paper’s abstract and main content to determine whether the paper was related to analytics techniques for MOOC videos or learning content. After collection and refining, we finally selected 74 representative papers for discussion. Fig. 2-a shows the research field of these papers. We categorized these papers by their research field, analytical task, employed data, and content of interest. Three co-authors collaboratively labeled each paper, each reviewing 1/3 of the papers and validating another 1/3 reviewed by the other co-authors. We referred to the categorization criteria from previous papers (Kui et al., 2022; Zhang et al., 2022), and a paper may have multiple labels. Discrepancies were resolved through multiple rounds of discussions with the three co-authors.
User Interview. Given that MOOC learning is primarily learner-driven (Lei et al., 2015; Blum-Smith et al., 2021), we interviewed 10 university students who had been using MOOCs for over three years (4 in computer science, 3 in finance, and 3 in art design majors). We aimed to understand learners’ behavioral patterns and concerns when watching MOOC videos, so we followed the findings from the literature to guide our interviews. All interviews were conducted in person. Participants were asked questions about MOOC learning, such as (1) When browsing a MOOC video, in what ways or viewing habits do they understand the course content, and (2) at what points during their MOOC video learning do they find it difficult to continue or feel the content is insufficient for their learning needs. Throughout the interview, they were free to browse courses on various MOOC platforms to facilitate their responses, and one co-author would record the learning process for further discussion.
Expert Feedback. Finally, we discussed the results of the literature review and user interviews with two experts in online learning (E1 & E2) to refine our design spaces and gather feedback. E1 is an instructor who has been offering online lecture videos on Coursera and edX platforms for 3 years, and E2 is an educational analyst who has a background in MOOC video research for 10 years. We conducted discussions with them face-to-face. We began by presenting the results of our literature review to establish the background of our work and potential categories for learning content. Next, we shared the results of our interviews with university students, focusing specifically on learners’ behavioral patterns while engaging with MOOC videos. Meanwhile, experts provided insights into the difficulties and popular analysis tasks they encountered in online learning. They also helped us refine the design space and suggested additional dimensions for analyzing potential cognitive aspects during the augmentation process. Through several rounds of discussions, we gained a better understanding of the design space and proposed a design space for augmenting MOOC videos based on expert feedback and insights gained from past augmented video design spaces(Chen et al., 2021).
@@ 3.2. Concept-based Design Space
Instructors integrate and organize concepts into videos, and learners complete learning by watching the entire video. This is the most basic delivery mode in online learning (Greenhow et al., 2022). Research in educational psychology and cognitive science indicates that comprehensive learning and understanding are usually achieved by constructing and connecting core concepts (Erickson, 2007; Latham, 2020). Our initial approach was to categorize all elements in MOOC videos to foster more general design practices. However, the literature review revealed that in various analytical tasks (Fig. 2-b), 79.7% (59 out of 74) of papers focus on concepts as the fundamental unit of analysis (Fig. 2-c), employing different representations of concepts (Fig. 2-d) to accomplish diverse analytical practices. Further discussions with experts showed that literature focusing on concept-based analysis aims to enhance learners’ understanding and progress in courses. Thus, embedding visualizations in MOOC videos must be concept-based, enabling learners to engage with the MOOC video thoroughly from the bottom up. This approach aligns well with MOOC video augmentation’s unique characteristics and requirements.
Compared to sport videos augmentation(Chen et al., 2021), MOOC videos are concept-driven, focusing on reinforcing learning outcomes through multimodal visualization and multi-stage interaction. They require complex navigation mechanisms to support diverse learning behaviors and emphasize conceptual connections. In contrast, sport videos augment viewers’ understanding of game elements within current frames. Inspired by this comparison, we propose a concept-based design space for augmenting MOOC videos, categorized into learning content-level and learner-level.

### v2

@@ 1. Introduction
With the rise of distance learning and online education platforms, Massive Open Online Courses (MOOCs) have continued to attract new learners over the past decade. However, learners often find it confusing to watch MOOC videos without the aid of additional knowledge(Milligan and Griffin, 2016) due to their lack of sufficient knowledge to understand the complex arrangement of the course and the diverse logical relationships of concepts - the fundamental ideas, theories, and principles that form the core of the course content. Existing methods of providing additional information, such as navigation menus (Zhang et al., 2019; Zhou et al., 2024) and course overviews (Rahman et al., 2023; Schwab et al., 2016), require learners to shift attention between different interfaces when explanatory materials or conceptual flows are placed outside the main video view. This increases cognitive load and disrupts conceptual continuity(Kizilcec et al., 2015). Thus, finding a way to seamlessly access additional data while watching a MOOC video is crucial for effectively learning and understanding a course.
@@
Embedded visualizations present a promising opportunity to access additional data by directly displaying them within actual scenes, enabling seamless transitions between real-time viewing and on-demand exploration without diverting user’s attention from MOOC videos. Various commercial products (Clippers, 2020; Vizrt, 2022) and research systems (Chen et al., 2023, 2021) have leveraged embedded visualizations to augment videos, but these systems either do not focus on the educational domain or only utilize basic non-interactive elements. Meanwhile, hypervideo (Chambel et al., 2006), as an emerging interactive video technology, allows users to access additional information on-demand by embedding hyperlinks within the video. Existing survey feedback indicates that hypervideo is regarded as an effective means in MOOCs for promoting knowledge acquisition and enhancing learner engagement (Milligan and Griffin, 2016). However, to the best of our knowledge, there is still no work focused on hypervideo-based augmentation for MOOC videos. MOOC videos are characterized by their short, modular format (often 5-15 minutes), high density of concepts, and the need to support diverse learners who learn asynchronously (Pilli and Admiraal, 2016). These distinct characteristics create a gap, as generic hypervideo solutions lack the specialized capabilities for automated, curriculum-aware augmentation that MOOC contexts require.
Understanding how to augment MOOC videos effectively requires examining both the content being taught and the learners engaging with it. Previous research has established that MOOC learning is fundamentally an interaction between two key components: the learning content (Object) and the learner (Subject) (Blum-Smith et al., 2021; Lei et al., 2015). This object-subject framework has been widely adopted in learning content analysis (Zawacki-Richter et al., 2018; Shih and Huang, 2005). However, augmenting MOOC videos within this framework presents unique challenges across both dimensions. At the learning content level, existing systems (Shih and Huang, 2005) categorize video content into multiple hierarchical levels, but provide limited guidance on how to effectively augment these different types of data. This raises two critical questions: What kinds of data can be used to augment MOOC videos (Q1)? What visual effects are used to augment these data (Q2)? At the learner level, MOOC learning is primarily learner-driven (Blum-Smith et al., 2021; Lei et al., 2015), where individual learners exhibit diverse viewing behaviors and cognitive states. Research on cognitive load theory has demonstrated that excessive cognitive load can significantly impair learning outcomes (Sweller et al., 1998; Chen et al., 2017). This introduces two additional challenges: How to dynamically augment data based on learners’ viewing behavior (Q3)? How to augment data without increasing the cognitive load on learners while promoting their understanding (Q4)?
@@
In this work, we aim to fill this gap by developing interactive embedded visualizations to help learners watch MOOC videos. Through systematic analysis of previous literature, learners’ feedback, and expert interviews, we derived a concept-based design space (Latham, 2020) to address the four aforementioned questions. The design space describes how to effectively augment MOOC videos in terms of both learning content and learner levels. The four dimensions of the design space (i.e., data type, visual effects, behavior level, and cognitive level) provide guidance for the augmentation of MOOC videos, offering new perspectives for facilitating learning in the online education domain.
@@
Guided by the design space, we designed and implemented HyperMOOC, a MOOC video augmentation tool that enhances learners’ engagement and understanding of MOOC videos Fig. 1-c). HyperMOOC uses raw MOOC videos as input (Fig. 1-a), extracts the elements at each element level (Sec. 3.3) through deep learning (DL) models, and implements a hyperlink-based interaction method to embed visualizations into MOOC video content (Fig. 1-b). Our interaction flow supports three stages—“Play”, “Focused”, and “Paused”—corresponding to continuous playback, focused attention on specific content, and pausing for deeper engagement. Using keyboard-mouse interaction, learners can seamlessly transition between these stages, balancing explanatory and exploratory aspects within MOOC video learning. Guided by the multi-stage methodology of knowledge acquisition (Mccarthy, 1996), we implemented concept-based glyphs to visually encode various types of basic data elements (figures & tables, equations & codes, examples & tests). A user study with 36 learners demonstrated the overall usability of HyperMOOC. Comparing three conditions (RAW, AUG, FULL), participants agreed that the FULL version can more effectively promote course understanding. We further collected feedback from two educational psychologists and discussed limitations and future research directions. The corpus, created videos, and other materials can be found in https://hypermooc.github.io/HyperMOOC/. In summary, the contributions of this paper can be summarized as follows:
@@ 2. Related Work
Object-level analysis focuses on indexing, retrieving, and browsing key concepts of online courses. Representative works include extracting index words from audio transcriptions(Haubold, 2004), applying OCR and keyword recognition to generate searchable term lists from lecture slides(Adcock et al., 2010), and unsupervised keyphrase extraction methods(Albahr et al., 2019), all reducing the difficulty of obtaining key concepts.
@@
Event-level analysis focuses on identifying patterns between concepts and analyzing their interrelationship during concept transitions. For instance, ConceptScape(Liu et al., 2018) identifies transitions such as algorithm $\rightarrow$ time complexity $\rightarrow$ optimization, capturing not only temporal order but also causal and hierarchical relationships between concepts. Concept mapping plays a crucial role in knowledge representation and assessment(Chang et al., 2001; Oliver, 2009). To reduce learners’ effort in crafting concept maps, tools like CmapTools(Cañas and Novak, 2014) and ScaffoMapping(Zhang et al., 2019) provide auxiliary functions to enhance mapping quality.
@@
Conclusion-level analysis focuses on providing course summaries to help learners grasp core concepts. Works in this area segment lecture videos with visual hints(Haubold and Kender, 2005), provide content overviews through keyframes(Zhao et al., 2016), and propose automated segmentation methods to improve navigation efficiency(Rahman et al., 2023).
Although the three levels of analysis offer complementary perspectives on learning content, existing systems typically treat them in isolation, lacking an integrated framework that connects object-, event-, and conclusion-level insights in a coherent design space, which limit learners’ ability to explore complex course materials holistically.
@@
Online Learning Video Visualizations. Numerous studies have explored diverse methods to visualize online learning videos, aiming to enhance educators’ and learners’ understanding and engagement with educational content, leading to understanding the value of online learning data from a visual analytics perspective(Emmons et al., 2017).
Visual analysis of clickstream data is crucial for tracking student progress. Systems such as PeakVizor(Chen et al., 2015), VUSphere(He et al., 2018), and ViSeq(Chen et al., 2018) employ various visualization techniques—clock-like glyphs, spherical layouts, and bidirectional chord diagrams—to analyze learner behaviors and correlate learning sequences with performance.
In terms of visualizing learning content, several works present concepts to make learning more engaging. Zhao et al.(Zhao et al., 2017) presented keyframe summaries in a comic book style. Torre et al.(Torre et al., 2022) proposed augmented transcripts and concept maps that dynamically highlight relevant segments, facilitating cognitive focus during video-based learning. Zhou et al.(Zhou et al., 2024) proposed ConceptThread, presenting course concepts in a thread view to facilitate understanding of concept relationships.
Existing work has yet to explore how visualizations can be embedded directly into MOOC videos to maintain learners’ contextual awareness without diverting attention away from the primary learning surface. To address this gap, our study aims to investigate how integrated video-embedded visualizations can support continuous situational awareness learning during video viewing.
@@
The dynamic nature of hypervideo transforms static information into engaging and motion-based content. For instance, Chambel et al.(Chambel et al., 2006) demonstrated that hypervideo could visualize static dimensions as dynamic ones, thus promoting comprehension. In MOOCs, this dynamism is particularly beneficial, allowing viewers to access additional information about video clips through hyperlinks.
Control features in hypervideo systems empower users to navigate and interact with video content more flexibly. HyperMeeting(Girgensohn et al., 2015) exemplifies this by offering extensive control over video playback, allowing participants to initiate playback for all current participants, navigate between previously meetings, and pause or skip recorded video streams.
Hyperlink is central to hypervideo, enabling the creation of interconnected video content and additional information. As seen in TrACE(Dorn et al., 2015), which allows viewers to annotate and embed discussions at specific locations and times in videos. Tiellet et al.(Tiellet et al., 2010) created MultiLinks, which allows viewing surgeries organized in steps, linked to corresponding times in the hypervideo, providing effective learning support in veterinary surgery education.
Although these systems demonstrate how navigation, annotation, and dynamic visualization can enrich educational video interactions, they typically treat these functions in isolation and rarely leverage them to uncover learners’ cognitive pathways or support conceptual understanding in MOOCs. HyperMOOC advances this paradigm by integrating conceptual visualizations, multi-symbol design, staged interactions, cognitive-path displays, and hyperlink-based navigation. This integrated design deepens learners’ understanding of MOOC content and knowledge structures while enhancing overall learning efficiency and experience.
@@ 3.1. Methodology
We employed a triangulated methodological approach combining literature review, user interviews, and expert consultations to establish a rigorous foundation for our design space, ensuring that our design choices are grounded in theoretical understanding, empirical evidence from learners, and practical insights from domain experts.
Literature Review. We conducted a systematic literature review using both search-driven and reference-driven methods. The search employed keywords “online learning”, “education videos”, “MOOC videos”, “learning content analysis”, and “e-learning analysis” across IEEE Xplore, ACM Digital Library, and Google Scholar, covering publications from 2010 to 2025. We focused on high-impact venues including conferences (IEEE VIS, ACM CHI, ACM L@S, ACM LAK) and journals (IEEE TVCG, Computer Graphics & Applications, Computer & Education). Following initial searches, we recursively examined references within identified papers. Papers were included if they focused on MOOC/online educational videos, presented analytical or visualization techniques, or investigated learner behaviors in video-based online education. This process yielded 74 representative papers (Fig. 2-a).
Three co-authors collaboratively coded these papers along four dimensions: research field, analytical task, employed data, and content of interest. Each author independently reviewed 1/3 of the papers while cross-validating another 1/3 reviewed by others, following categorization criteria from prior work (Kui et al., 2022; Zhang et al., 2022). Papers could receive multiple labels when applicable, and discrepancies were resolved through iterative discussions until consensus was reached.
User Interview. Given that MOOC learning is primarily learner-driven (Lei et al., 2015; Blum-Smith et al., 2021), we conducted semi-structured interviews with 10 university students (4 computer science, 3 finance, 3 art design majors) who had used MOOCs actively for over 3 years and completed at least 5 courses on major platforms. Participants were recruited through university mailing lists. The 45-60 minute face-to-face interviews investigated: (1) how participants approach and process content when watching MOOC videos, including viewing habits and comprehension tactics, and (2) specific points where they encounter difficulties or require additional support. The interview protocol was developed based on findings from our literature review to ensure alignment with documented challenges.
During interviews, participants browsed actual MOOC platforms to demonstrate typical learning behaviors and articulate their thought processes in real-time. One co-author documented each session, recording both verbal responses and observed behaviors. Interview data were analyzed using iterative coding to identify recurring patterns and learner needs that informed our design requirements.
Expert Consultation. We conducted structured consultation sessions with two domain experts who bring complementary perspectives on MOOC pedagogy and analytics. E1 is a MOOC instructor with 3 years of experience developing online lecture videos on Coursera and edX, providing practical insights from teaching thousands of students. E2 is an educational researcher with 10 years of experience in MOOC analytics, offering theoretical grounding and empirical evidence from large-scale learning analytics studies.
The face-to-face consultation sessions (2̃ hours each) followed a systematic protocol: First, we presented our literature review findings to establish shared understanding of the research landscape. Second, we shared synthesized results from student interviews, focusing on learner behavioral patterns and identified pain points. Third, we solicited expert perspectives on common difficulties in online learning, evaluation of our proposed design dimensions, and refinements based on their domain knowledge. This expert consultation served as critical triangulation, validating that our design choices integrated theoretical foundations, empirical evidence from learners, and practical wisdom from experienced practitioners and researchers.
@@ 3.2. Concept-based Design Space
Instructors integrate and organize concepts into videos, and learners complete learning by watching the entire video. This is the most basic delivery mode in online learning (Greenhow et al., 2022). The necessity of concept-based augmentation is grounded in established theories of learning and cognition. Ausubel’s meaningful learning theory (Ausubel et al., 1978) posits that “the most important single factor influencing learning is what the learner already knows,” emphasizing that learners construct understanding by relating new information to existing conceptual frameworks. Building on this foundation, Erickson’s concept-based learning model (Erickson, 2007) demonstrates that concepts, as abstract, timeless, and universal ideas, serve as the cognitive bridge enabling knowledge transfer across contexts, a critical requirement for effective MOOC learning (Latham, 2020; Lasater and Nielsen, 2009). Unlike isolated facts that remain “locked in time and place,” concepts facilitate deeper understanding through term “synergistic thinking”, the cognitive interplay between factual and conceptual levels of mental processing (Erickson, 2007).
This theoretical foundation is empirically supported by our literature review, which revealed that 79.7% (59 out of 74) of papers analyzing MOOC videos focus on concepts as the fundamental unit of analysis (Fig. 2-b,c), employing different representations of concepts (Fig. 2-d) to accomplish diverse analytical tasks. Further discussions with domain experts confirmed that concept-based analysis consistently leads to enhanced learner understanding and course progress. The pedagogical benefits are substantial: concept-based approaches promote transferable understanding rather than rote memorization, enable learners to see connections between disparate topics, and support the construction of coherent mental models, all essential for self-directed MOOC learning where instructional scaffolding is limited (Erickson, 2007; Latham, 2020; Lasater and Nielsen, 2009).
Thus, embedding visualizations in MOOC videos must be concept-based, enabling learners to engage with the MOOC video thoroughly from the bottom up. This approach aligns well with MOOC video augmentation’s unique characteristics and requirements.